# MLC at Feigenbaum points

### v2

@@ 3.5.8. Almost periodic rectangles
An almost periodic rectangle $R$ between little bushes ${\mathfrak{B}}^{[n]}_{a},{\mathfrak{B}}^{[n]}_{a+1}$ is defined in the same way as in the case $n=1$. Such a rectangle $R$ can be lifted to the dynamical plane of $F_{n-1,c}$ (via the covering map representing the $\psi^{\bullet}$-renormalization, see §3.4.3) and its lift will be an almost periodic rectangle between level 11 11 little bushes as in Definition 3.12 .
@@ 4. Pull-off for non-periodic rectangles
Consider the local part ${\mathcal{A}}^{k}_{i}$ of ${\mathcal{A}}^{k}$ around ${\mathfrak{B}}^{n}_{i}$. Write $q\coloneqq(10\overline{p}+3)p_{n}$. If for all $i$ we have
@@
${\mathcal{W}}\left(V\setminus{\mathfrak{K}}_{j}^{[n]}\right)\asymp_{\delta_{1},\overline{p}}\ K\hskip 14.22636pt\text{ for all }j.$
@@
${\mathcal{W}}(F)\geq\sum_{j}{\mathcal{W}}\left(V\setminus{\mathfrak{K}}_{j}^{[n]}\right)\succeq_{\delta_{1},\overline{p}}\ p_{n}K\gg 2K,$
@@ 4.1. Alignment of WAD with the Hubbard continuum
(4.2) ${\mathcal{A}}^{(m+1),k+1}_{{\mathrm{hor}}}=f^{*}\left({\mathcal{A}}^{(m),k}_{{\mathrm{hor}}}\right){\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\mathcal{A}}^{(m),k+1}_{\mathrm{hor}}\geq{\mathcal{A}}^{(m),k}_{\mathrm{hor}}.$
Let us choose a sufficiently big threshold $C\gg_{p_{n}}1$. By (4.2), the ADs forming ${\mathcal{A}}^{k}_{\mathrm{hor}}-C^{k}$ decrease, hence for a certain
@@
Since the WAD ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}$ is dominated by the WAD $f^{*}\left({\mathcal{H}}\right)=f^{*}\left({\mathcal{H}}-{\mathcal{A}}^{\mathbf{t}}_{\mathrm{hor}}-C^{\mathbf{t}}\right)$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}^{(1)}$, see §3.5.7, we obtain that $H$ is an invariant AD, and therefore it is aligned with the Hubbard continuum by Lemma 2.2.
@@ 4.2. Case (P)
[…]
@@
${\mathcal{W}}^{\bullet}(F)={\mathcal{W}}(V,\ {\mathfrak{B}}_{F})\geq{\mathcal{W}}(V,\ {\boldsymbol{\mathfrak{B}}})\geq 2K.$
@@ 4.3. Case (S)
[…]
@@
Write $q^{\prime}=10\overline{p}p_{n}$. Let ${\mathcal{H}}^{(q^{\prime})}_{c}$ be the WAD consisting of arcs of $\big{(}f^{q^{\prime}}\big{)}^{*}{\mathcal{H}}$ that are in ${\mathfrak{B}}^{[n-1]}_{c}$. By Lemma 3.11, the restriction $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ consists of horizontal curves. Since $H={\operatorname{AD}}({\mathcal{H}})$ is aligned with the Hubbard dendrite, $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ is, after removing $O_{p_{n}}(1)$ curves, dominated by ${\mathcal{H}}^{(q^{\prime})}_{c}$, see §3.5.7. And since $(f^{q^{\prime}})_{*}\colon{\mathcal{H}}^{q^{\prime}}_{c}\to{\mathcal{H}}_{c}$ is a bijection, as most $\overline{p}\delta_{1}K+2C^{2}<\delta_{2}K$ curves in ${\mathcal{H}}_{c}$ can have expansivity number greater than 22 22, where $\delta_{1}\ll\delta_{2}\ll\delta$. This implies that most curves in ${\mathcal{F}}({\mathcal{H}}_{c})$ are within rectangles representing genuine periodic arcs §2.4.3 that connect neighboring level $n$ bushes of ${\mathfrak{B}}^{[n-1]}_{c}$, see Lemma 2.3. This also demonstrates that all the ${\mathcal{H}}_{c}$ have comparable width – the difference of their weights are bounded by $\delta_{2}K$.
@@ 5. Waves
Given a compact connected filled set $X\subset{\mathbb{C}}$, we denote by $\partial^{c}X$ its Carathéodory boundary; i.e., the set of prime ends of ^$\widehat{\mathbb{C}}\setminus X$. A Riemann map identifies $\partial^{c}\widetilde{X}$ with the unit circle ${\mathbb{T}}$. For a compact connected subset $Y\subset X$, we denote $\partial^{c}_{X}Y\subset\partial^{c}X$ the set of prime ends of $X$ accumulating at $Y$. A side of $Y$ rel $X$ is a connected component of $\partial^{c}_{X}Y$ viewed as a subset of ${\mathbb{T}}\simeq\partial^{c}X$.
@@
  • curves in ${\mathcal{S}}^{t}$, $t>0$ start at an interval $I^{t}_{-}\subset A_{1}$ and end at an interval $I^{t}_{+}\subset A_{1}$ such that one of the intervals $I^{t}_{-},I^{t}_{+}$ is within a union of $1\displaystyle 2^{t}-$ 1 fundamental intervals for the action $H\curvearrowright A_{1}$ and the interval $J^{t}\subset A_{1},$ $J^{t}\cap J\not=\emptyset$ between $I^{t}_{-},I^{t}_{+}$ is the union of exactly $1\displaystyle 2^{t}-$ 1 fundamental intervals of the action $H\curvearrowright A_{1}$;
@@
Then ${\mathcal{W}}({\mathcal{S}}^{t})=O(1)$ for $t>0$ because ${\mathcal{S}}^{t}$ crosses its shift under $H$, see [DL2, Shift Argument §A.3].
Replacing ${\mathcal{S}}$ with ${\mathcal{S}}^{0}$, we obtain that curves in the ${\mathcal{S}}^{0}_{k}$ start and end at pairwise disjoint intervals. Removing an extra $0(1)$ outermost curves, we obtain that the new ${\mathcal{S}}^{\operatorname{new}}_{l}$ are pairwise disjoint. ∎
Let us assume first that ${\mathfrak{K}}_{k}$ is periodic. Fix an iteration $f^{n}\mid T,\ n\leq M^{3}$ such that $f^{n}\mid{\mathfrak{K}}_{k}^{\iota}$ covers $\partial^{c}_{T}{\mathfrak{K}}_{k}$ at least twice. Consider the associated iteration
@@
Let ${\mathcal{F}}^{\prime}$ be the family of all $\gamma$ in ${\mathcal{S}}$ with $\gamma_{a}\not=\gamma_{b}$. Let ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ be family of curves consisting of $\gamma_{a}$ and $\gamma_{b}$ with $\gamma$ in ${\mathcal{S}}^{\prime}$. Since ${\mathcal{S}}^{\prime}$ overflows ${\mathcal{S}}_{a}$ and then ${\mathcal{S}}_{b}$, either ${\mathcal{W}}({\mathcal{S}}_{a})$ or ${\mathcal{W}}({\mathcal{S}}_{b})$ is at least $4/3K-O_{M}(1)$. After removing $O_{M}(1)$ buffers and applying Lemma 5.2, the waves ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ map univalently by onto waves above ${\mathfrak{K}}_{k}^{\iota}$. We obtain a wave with width $\geq\frac{4}{3}K-O_{M}(1)$ and the whole argument can be repeated with a strictly bigger wave leading eventually to the contradiction.
@@ 6.1.1. Spiraling Numbers
${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(q)}_{a}$${\mathfrak{B}}^{\prime\prime}_{1}$$\ell_{1}$${\mathfrak{B}}^{\prime}_{1}$${\mathfrak{B}}_{i}^{[2]}$${\mathfrak{B}}_{i+1}^{[2]}$${\mathfrak{B}}^{\prime\prime}_{2}$${\mathfrak{B}}^{\prime}_{2}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{a+1}$${\mathfrak{B}}^{[1],(q)}_{a+1}$${\mathcal{G}}_{a}$$\alpha$$R^{(q)}$
Figure 11. The non-spiraling case. The rectangle $R$ (orange, violet, red) connects ${\mathfrak{B}}^{[1]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a+1}$. Between ${\mathfrak{B}}^{[1],(q)}_{a}$ and ${\mathfrak{B}}^{[1],(q)}_{a+1}$ most of the rectangle $R$ travels through its lift $R^{(q)}$ (red). Therefore, the part ${\mathcal{G}}_{a}$ (violet) of $R$ between $\ell_{1}$ and $\ell_{2}$ is tremendously wide. The bushes ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1},{\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ are lifts of ${\mathfrak{B}}^{[1]}_{a+1}$, see Figure 6.
@@
Recall the the pure Mapping Class Group $\operatorname{MCG}$ of a 33 33-punctured sphere $S^{2}\setminus\{a,b,\infty\}$ is trivial. By replacing $a,b$ with small open Jordan disks $D_{a},D_{b}$, we obtain the new group $\operatorname{MCG}(S^{2}\setminus(D_{a}\cup D_{b}\cup\{\infty\}))\simeq{\mathbb{Z}}^{2}$ consisting of Dehn-twists around $D_{a},D_{b}$, see [FM]. This fact has the following implication.
Let $O$ be a small disk-neighborhood of ${\mathfrak{B}}^{[1]}_{a}\cup\Pi\cup{\mathfrak{B}}^{[1]}_{a+1}$. Then $O^{\prime}\coloneqq O\setminus({\mathfrak{B}}^{[1]}_{a}\cup{\mathfrak{B}}^{[1]}_{a+1})$ is topologically a disk with two holes. Therefore, a path (simple curve) $\beta\subset O^{\prime}$ from ${\mathfrak{B}}^{[1]}_{a}$ to ${\mathfrak{B}}^{[1]}_{a+1}$ has a simple description: it first spirals $\tilde{t}_{a}\in{\mathbb{Z}}$ times around ${\mathfrak{B}}^{[1]}_{a}$ and then spirals $\tilde{t}_{a+1}\in{\mathbb{Z}}$ times around ${\mathfrak{B}}^{[1]}_{a+1}$, where $\Pi$ can be used as a reference “path” of zero spiralings. Below we will ignore the spiraling orientation and introduce the absolute quantity $t_{a}=|\tilde{t}_{a}|$.
@@
  • $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a}$ for $t\leq t_{a}$;
  • $\gamma_{t}\cup\Pi$ does not surround ${\mathfrak{B}}^{[1]}_{a}$ for $t>t_{a}$.
@@
    The spiraling numbers of $\gamma_{1},\gamma_{2}\in R$ differ by at most 11 11.
@@
    If $\gamma_{s}\in R^{(s)}$ is the lift of $\gamma\in R$ into $R^{(s)}$, then $t_{a}(\gamma^{s})\leq\frac{1}{2^{s}}t_{a}(\gamma)$.
@@ 6.1.2. The non-spiraling case:
There are two possibilities. If a substantial part of ${\mathcal{G}}_{a}$ travels through both ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$, then this substantial part of ${\mathcal{G}}_{a}$ restricts to a wide lamination ${\mathcal{L}}$ between ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$. Pushing ${\mathcal{L}}$ with respect to $f_{*}$ and $\iota^{*}$ and using Lemma 3.2 (the number of iterations and the degree is bounded in terms of $\overline{p}$), we obtain a family ${\mathcal{F}}$ of non-trivial proper curves in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[2]}$ starting at ${\mathfrak{B}}^{[2]}_{0}$ such that ${\mathcal{W}}({\mathcal{F}})\succeq_{\overline{p}}C_{0,\delta}K$. If a substantial part of ${\mathcal{F}}$ is vertical, then we obtain the second estimate in (6.1), (6.2). If a substantial part of ${\mathcal{F}}$ is horizontal, then we obtain the first estimate in (6.1); applying Lemma 3.10 to ${\mathcal{F}}$, we obtain the first estimate in (6.2).
If a substantial part of ${\mathcal{G}}_{a}$ omits either ${\mathfrak{B}}^{[2]}_{i}$ or ${\mathfrak{B}}^{[2]}_{i+1}$, then we obtain a wide wave above one of the sides of ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$; Wave Lemma 5.1 implies that ${\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K$.
[…]

### v3

@@ 3.5.8. Almost periodic rectangles
An almost periodic rectangle $R$ between little bushes ${\mathfrak{B}}^{[n]}_{a},{\mathfrak{B}}^{[n]}_{a+1}$ is defined in the same way as in the case $n=1$. Such a rectangle $R$ can be lifted to the dynamical plane of $F_{n-1,c}$ (via the covering map representing the $\psi^{\bullet}$-renormalization, see §3.4.3) and its lift will be an almost periodic rectangle between level $1$ little bushes as in Definition 3.12 .
@@ 4. Pull-off for non-periodic rectangles
Consider the local part ${\mathcal{A}}^{k}_{i}$ of ${\mathcal{A}}^{k}$ around ${\mathfrak{B}}^{n}_{i}$. Write $q\coloneqq(10\overline{p}+3)p_{n}$. We remark that since $q$ is linear in $p_{n}$, the $q$th iterate has a bounded degree on small bushes and the below application of the Covering Lemma will lead to estimates independent of $n$.
If for all $i$ we have
@@
${\mathcal{W}}\left(V\setminus{\mathfrak{B}}_{j}^{[n]}\right)\asymp_{\delta_{1},\overline{p}}\ K\hskip 14.22636pt\text{ for all }j.$
@@
${\mathcal{W}}(F)\succeq_{\delta_{1},\overline{p}}\ \sum_{j}{\mathcal{W}}\left(V\setminus{\mathfrak{B}}_{j}^{[n]}\right)\succeq_{\delta_{1},\overline{p}}\ p_{n}K\gg 2K,$
@@ 4.1. Alignment of WAD with the Hubbard continuum
(4.2) ${\mathcal{A}}^{(m+1),k+1}_{{\mathrm{hor}}}=f^{*}\left({\mathcal{A}}^{(m),k}_{{\mathrm{hor}}}\right){\ \ }{\ \ }\text{ and }{\ \ }{\ \ }{\mathcal{A}}^{(m),k+1}_{\mathrm{hor}}\leq{\mathcal{A}}^{(m),k}_{\mathrm{hor}}.$
Let us choose a sufficiently big threshold $C\gg_{p_{n}}1$ so that (3.29) is applicable for $m=0$. By (4.2), the ADs forming ${\mathcal{A}}^{k}_{\mathrm{hor}}-C^{k}$ decrease, hence for a certain
@@
Since the WAD ${\mathcal{A}}^{{\mathbf{t}}+1}_{\mathrm{hor}}-C^{{\mathbf{t}}+1}$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}$ is dominated by the WAD $f^{*}\left({\mathcal{H}}\right)=f^{*}\left({\mathcal{A}}^{\mathbf{t}}_{\mathrm{hor}}-C^{\mathbf{t}}\right)$ on $U^{{\mathbf{t}}+1}\setminus{\boldsymbol{\mathfrak{B}}}^{(1)}$, see (3.29) (where $M=C^{\mathbf{t}}$) in §3.5.7, we obtain that $H$ is an invariant AD, and therefore it is aligned with the Hubbard continuum by Lemma $2.$
@@ 4.2. Case (P)
Applying [KL1, Quasi Additivity Law with separation] together with (4.6) and using $p_{n}\gg_{\delta_{1},\overline{p}}\ 1$, we obtain
@@
${\mathcal{W}}_{\bullet}(F)={\mathcal{W}}(V,\ {\mathfrak{B}}_{F})\geq{\mathcal{W}}(V,\ {\boldsymbol{\mathfrak{B}}})\geq 2K.$
This establishes Case (P) from (Ploc).
@@ 4.3. Case (S)
Let us now assume that (Ploc) does not hold: for all $i$ we have ${\mathcal{W}}({\mathcal{A}}_{i,{\mathrm{ver}}}^{q})\leq\delta_{1}K$. This implies that most paths in the canonical lamination ${\mathcal{F}}({\mathcal{A}}_{i})$ of ${\mathcal{A}}_{i}$ are horizontal and they restrict to horizontal paths under the immersions $\iota^{s}\colon(U^{s},{\boldsymbol{\mathfrak{B}}})\to(V,{\boldsymbol{\mathfrak{B}}})$ for $s\leq q$. We will use notations of §2.4.4.
@@
Write $q^{\prime}=10\overline{p}p_{n}$. Let ${\mathcal{H}}^{(q^{\prime})}_{c}$ be the WAD consisting of arcs of $\big(f^{q^{\prime}}\big)^{*}{\mathcal{H}}$ that are in ${\mathfrak{B}}^{[n-1]}_{c}$. By Lemma 3.11, the restriction $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ consists of horizontal curves. Since $H={\operatorname{AD}}({\mathcal{H}})$ is aligned with the Hubbard dendrite, $(\iota^{q^{\prime}})^{*}\left[{\mathcal{H}}_{c}-\delta_{1}K\right]$ is, after removing $O_{p_{n}}(1)$ curves, dominated by ${\mathcal{H}}^{(q^{\prime})}_{c}$, see §3.5.7. And since $(f^{q^{\prime}})_{*}\colon{\mathcal{H}}^{q^{\prime}}_{c}\to{\mathcal{H}}_{c}$ is a bijection, as most $\overline{p}\delta_{1}K+2C^{2}<\delta_{2}K$ curves in ${\mathcal{H}}_{c}$ can have expansivity number greater than $2$, where $\delta_{1}\ll\delta_{2}\ll\delta$. This implies that most curves in ${\mathcal{F}}({\mathcal{H}}_{c})$ are within rectangles representing genuine periodic arcs §2.4.3 that connect neighboring level $n$ bushes of ${\mathfrak{B}}^{[n-1]}_{c}$, see Lemma 2.3. This also demonstrates that all the ${\mathcal{H}}_{c}$ have comparable width – the difference of their weights are bounded by $\delta_{2}K$.
@@ 5. Waves
Given a compact connected filled set $X\subset{\mathbb{C}}$, we denote by $\partial^{c}X$ its Carathéodory boundary; i.e., the set of prime ends of $\widehat{\mathbb{C}}\setminus X$. A Riemann map identifies $\partial^{c}\widetilde{X}$ with the unit circle ${\mathbb{T}}$. For a compact connected subset $Y\subset X$, we denote $\partial^{c}_{X}Y\subset\partial^{c}X$ the set of prime ends of $X$ accumulating at $Y$. A side of $Y$ rel $X$ is a connected component of $\partial^{c}_{X}Y$ viewed as a subset of ${\mathbb{T}}\simeq\partial^{c}X$.
@@
  • curves in ${\mathcal{S}}^{t}$, $t>0$ start at an interval $I^{t}_{-}\subset A_{1}$ and end at an interval $I^{t}_{+}\subset A_{1}$ such that one of the intervals $I^{t}_{-},I^{t}_{+}$ is within a union of $2^{t}-1$ fundamental intervals for the action $H\curvearrowright A_{1}$ and the interval $J^{t}\subset A_{1},$ $J^{t}\cap J\not$$\emptyset$ between $I^{t}_{-},I^{t}_{+}$ is the union of exactly $2^{t}-1$ fundamental intervals of the action $H\curvearrowright A_{1}$;
@@
Then ${\mathcal{W}}({\mathcal{S}}^{t})=O(1)$ for $t>0$ because, after removing $1$ buffers, ${\mathcal{S}}^{t}$ crosses its shift under $H$, see [DL2, Shift Argument §A.3].
Replacing ${\mathcal{S}}$ with ${\mathcal{S}}^{0}$, we obtain that curves in the ${\mathcal{S}}^{0}_{k}$ start and end at pairwise disjoint intervals. Removing an extra $O(1)$ outermost curves, we obtain that the new ${\mathcal{S}}^{\operatorname{new}}_{l}$ are pairwise disjoint. ∎
Let us assume first that ${\mathfrak{K}}_{k}$ is periodic. Since the number sides is $M$ and since the period of small Julia sets is bounded by $M$, we can fix an iteration $f^{n}\mid T$, where $n$ is bounded in terms of $M$ (for instance, $\ n\leq M^{3}$) such that $f^{n}\mid{\mathfrak{K}}_{k}^{\iota}$ covers $\partial^{c}_{T}{\mathfrak{K}}_{k}$ at least twice. Consider the associated iteration
@@
Write $K\coloneqq{\mathcal{W}}({\mathcal{S}})$. Let ${\mathcal{F}}^{\prime}$ be the family of all $\gamma$ in ${\mathcal{S}}$ with $\gamma_{a}\not=\gamma_{b}$. Let ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ be family of curves consisting of $\gamma_{a}$ and $\gamma_{b}$ with $\gamma$ in ${\mathcal{S}}^{\prime}$. Since ${\mathcal{S}}^{\prime}$ overflows ${\mathcal{S}}_{a}$ and then ${\mathcal{S}}_{b}$, either ${\mathcal{W}}({\mathcal{S}}_{a})$ or ${\mathcal{W}}({\mathcal{S}}_{b})$ is at least $4/3K-O_{M}(1)$. After removing $O_{M}(1)$ buffers and applying Lemma 5.2, the waves ${\mathcal{S}}_{a},{\mathcal{S}}_{b}$ map univalently by $f^{n}$ onto waves above ${\mathfrak{K}}_{k}^{\iota}$. We obtain a wave with width $\geq\frac{4}{3}K-O_{M}(1)$ and the whole argument can be repeated with a bigger wave leading eventually to the contradiction.
@@ 6.1.1. Spiraling Numbers
[…]
Figure 11. The non-spiraling case. The rectangle $R$ (orange, violet, red) connects ${\mathfrak{B}}^{[1]}_{a}$ and ${\mathfrak{B}}^{[1]}_{a+1}$. Between ${\mathfrak{B}}^{[1],(q)}_{a}$ and ${\mathfrak{B}}^{[1],(q)}_{a+1}$ most of the rectangle $R$ travels through its lift $R^{(q)}$ (red). Therefore, the part ${\mathcal{G}}_{a}$ (violet) of $R$ between $\ell_{1}$ and $\ell_{2}$ is tremendously wide. The bushes ${\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime\prime\prime},{\mathfrak{B}}^{\text{iv}}$ are lifts of ${\mathfrak{B}}^{[1]}_{a+1}$, see Figure 6.
@@
Recall the the pure Mapping Class Group $\operatorname{MCG}$ of a $3$-punctured sphere $S^{2}\setminus\{a,b,\infty\}$ is trivial. By replacing $a,b$ with small open Jordan disks $D_{a},D_{b}$, we obtain the new group $\operatorname{MCG}(S^{2}\setminus(D_{a}\cup D_{b}\cup\{\infty\}))\simeq{\mathbb{Z}}^{2}$ consisting of Dehn-twists around $D_{a},D_{b}$, see [FM]. This fact has the following implication.
[…]
@@
  • $\gamma_{t}\cup\Pi$ surrounds ${\mathfrak{B}}^{[1]}_{a}$ for $t\in\{1,2,\dots,t_{a}\}$;
  • $\gamma_{t}\cup\Pi$ does not surround ${\mathfrak{B}}^{[1]}_{a}$ for $t\in\{t_{a}+1,\dots,f\}$.
@@
    The spiraling numbers of $\gamma_{1},\gamma_{2}\in R$ differ by at most $1$.
@@
    If $\gamma_{s}\in R^{(s)}$ is the lift of $\gamma\in R$ into $R^{(s)}$, then $t_{a}(\gamma_{s})\leq\frac{1}{2^{s}}t_{a}(\gamma)$.
@@ 6.1.2. The non-spiraling case:
There are two possibilities. If a substantial part of ${\mathcal{G}}_{a}$ travels through both ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$, then this substantial part of ${\mathcal{G}}_{a}$ restricts to a wide lamination ${\mathcal{L}}$ between ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$. Pushing ${\mathcal{L}}$ with respect to $f_{*}$ and $\iota^{*}$ and using Lemma 3.2 (the number of iterations and the degree is bounded in terms of $\overline{p}$), we obtain a family ${\mathcal{F}}$ of non-trivial proper curves in $V\setminus{\boldsymbol{\mathfrak{B}}}^{[2]}$ starting at ${\mathfrak{B}}^{[2]}_{0}$ such that ${\mathcal{W}}({\mathcal{F}})\succeq_{\overline{p}}C_{0,\delta}K$. If a substantial part of ${\mathcal{F}}$ is vertical, then we obtain the second estimate in (6.1), (6.2). If a substantial part of ${\mathcal{F}}$ is horizontal, then we obtain the first estimate in (6.1); applying Lemma 3.10 to ${\mathcal{F}}$, we obtain the first estimate in (6.2).
If a substantial part of ${\mathcal{G}}_{a}$ omits either ${\mathfrak{B}}^{[2]}_{i_{1}}$ or ${\mathfrak{B}}^{[2]}_{i_{2}}$, then we obtain a wide wave above one of the sides of ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$; Wave Lemma 5.1 implies that ${\mathcal{W}}_{\bullet}(F)\geq C_{\delta}K$.
${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}_{a}^{[1],(1)}$${\mathfrak{B}}^{[1]}_{a+1}$${\mathfrak{B}}_{a+1}^{[1],(1)}$${\mathcal{G}}_{a}$$R^{(1)}$$\alpha$